# Efficient designs for multivariate crossover trials

### v1

@@ 3.2.1 Study of design efficiency for the markovian covariance structure
$\bm{V}_{R}^{*}$ $=\bm{V}_{R}^{-1}-\left(\bm{{1}}_{p}^{{}^{\prime}}\bm{V}_{R}^{-1}\bm{{1}}_{p}\right)^{-1}\bm{V}_{R}^{-1}\bm{J}_{p\times p}\bm{V}_{R}^{-1}=\bm{H}_{3},$
where $\bm{V}_{1}=\sigma_{11}\begin{bmatrix}1&r_{(1)}&r_{(1)}^{2}\\ r_{(1)}&1&r_{(1)}\\ r_{(1)}^{2}&r_{(1)}&1\end{bmatrix}$ and $\bm{V}_{R}=\bm{I}_{p}$. Here, $\sigma_{11}=\frac{\gamma_{11}}{1-r_{(1)}^{2}}$, $\gamma_{11}>0$ and $10<|r_{(1)}|<$ 1. Note $r_{(1)}=0$ reduces the $AR(1)$ structure to the diagonal structure, so we take $r_{(1)}\neq 0$. Also, $|r_{(1)}|<1$ is required for the matrix $\bm{V}_{1}$ to be positive definite. To check the necessary and sufficient condition from Remark 3.3, we evaluate $tr\left(\bm{V}_{1}^{*}\bm{\psi}\right)=-\gamma_{11}^{-1}\frac{2+r_{(1)}}{3-r_{(1)}}$, $tr\left(\bm{V}_{R}^{*}\bm{\psi}\right)=-\frac{2}{3}$, $tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{1}^{*}\bm{\psi}\right)=\gamma_{11}^{-1}\frac{10+6r_{(1)}}{3\left(3-r_{(1)}\right)}$ and $tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)=\frac{10}{9}$. Since $tr\left(\bm{V}_{1}^{*}\bm{\psi}\right)tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)$ and $tr\left(\bm{V}_{R}^{*}\bm{\psi}\right)tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{1}^{*}\bm{\psi}\right)$ are unequal, we may conclude that for $10<|r_{(1)}|<$ 1, $d^{*}$ does not attain the upper bound of $tr\left(\bm{{C}}_{d(s2)}\right)$ over $d\in\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$. We find that similar results hold for the other cases when $p=t=3$, and for all the cases when $p=t=4$. Thus we may state that for $p=t\in\{3,4\}$, and all cases of $\bm{V}_{1}$ and $\bm{V}_{R}$ considered, $d^{*}$ fails to attain the upper bound $u$.
@@
Here, $RD_{d^{(0)}}$ is a function of $t$, $n$, $p$, $\sigma_{11}$, $\sigma_{22}$, $\rho$, $\bm{V}_{1}$ and $\bm{V}_{R}$, and its values lie between 00 and $1$. For $RD_{d^{(0)}}=0$, the upper bound is attained by the design $d^{(0)}$, thus proving $d^{(0)}$ is an efficient design. Thus lower values of $RD_{d^{(0)}}$ are preferred, as they indicate that $tr\left(\bm{{C}}_{d^{(0)}(s2)}\right)$ is closer to the upper bound $u$, and $d^{(0)}$ is a highly efficient design.
@@
The balanced uniform design $d_{2}$ is constructed using the method described by Bose and Dey, (2009). For $p=t=3$, a balanced uniform design with the same number of subjects as $d^{*}$ is also an orthogonal array, so it is not considered as a candidate for comparison with $d^{*}$. A graphical procedure is used to evaluate the designs $d^{(0)}\in\{d_{1},d_{2},d^{*}\}$. In our illustrations we assume $\sigma_{11}=\sigma_{22}$. The next remark, shows that if $\sigma_{11}=\sigma_{22}$, then for $p=t\geq 3$, the relative difference does not depend on the value of $\sigma_{11}$.
@@
From our choices of $\bm{V}_{1}$ and $\bm{V}_{R}$, it is clear that $\bm{V}_{1}$ is a function of $\sigma_{11}$ and $r_{(1)}$, while $\bm{V}_{R}$ is only a function of $r_{(1)}$. In $\bm{V}_{1}$, $r_{(1)}$ is the entry just above/below the main diagonal in the correlation matrix corresponding to $\bm{V}_{1}$, while in $\bm{V}_{R}$, $r_{(1)}$ is the entry just above/below the main diagonal. So, these two matrices are diagonal if and only if $r_{(1)}=0$. Note, we have considered that when $\bm{V}_{1}$ is $AR(1)$ or tridiagonal, $\bm{V}_{R}$ is diagonal and when $\bm{V}_{1}$ is diagonal, $\bm{V}_{R}$ is $AR(1)$ or tridiagonal. Hence, in each case, only one of $\bm{V}_{1}$ or $\bm{V}_{R}$ is a function of non-zero $r_{(1)}$. Thus from Remark 3.4, we can see that for any $d^{(0)}\in\mathcal{D}^{(1)}_{t,n=\lambda t\left(t-1\right),p=t}$, $p=t\geq 3$ and $\sigma_{11}=\sigma_{22}$, the relative difference given as $RD_{d^{(0)}}$ is a function of $d^{(0)}$, $p$, $n$, $\rho$ and $r_{(1)}$.
@@
For $p=t\in\{3,4\}$, $n=t\left(t-1\right)$, $\sigma_{11}=\sigma_{22}$, and $d^{(0)}\in\{d_{1},d_{2},d^{*}\}$, we plot $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ and $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$, corresponding to various non-zero permissible values of $r_{(1)}$. Here, by permissible values, we mean those values for which the matrices $\bm{V}_{1}$ and $\bm{V}_{R}$ are positive definite. If in any case $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ attains 00, then this implies that design $d^{(0)}$ maximizes $tr\left(\bm{{C}}_{d^{(0)}(s2)}\right)$ and $d^{(0)}$ is a efficient design. If $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ does not attain 00, we plot $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ to get an understanding of the worst scenario and quantify the departure of the design under consideration from an efficient design. Lower the values of $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$, better is the performance of the design in terms of efficiency. For $d^{(0)}\in\{d_{1},d^{*}\}$, where $p=t=3$ and $n=6$, Figure 2 displays the plots of $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ and $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$, for all cases. While, for $p=t=4$, $n=12$, Figure 3 illustrates the performance of the three designs.
@@
From Figure 2, for $p=t=3$ and $n=6$, we see that $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d_{1}}$ never attains 00 and hence $d_{1}$ like $d^{*}$ fails to attain the upper bound $u$. Also, we note that for all cases the values of $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d_{1}}>>max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{*}}$, implying that $d^{*}$ performs much better when compared to $d_{1}$. Considering all four cases, the maximum value of $RD_{d^{*}}$ over $\rho\neq 0$ is at most around $4\%$ indicating that $d^{*}$ is a highly efficient design. In the $p=t=4$ and $n=12$ case, we observe from Figure 3 that none of the designs $d^{(0)}\in\{d_{1},d_{2},d^{*}\}$, $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ attains 00. In Figure 3, for all the cases, we see that $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{*}}<max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d_{2}}<max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d_{1}}$. So, $d^{*}$ performs the best among the three designs being compared, followed by $d_{2}$. Since $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d_{1}}$ is far away from 00, we reject the candidature of $d_{1}$ as a highly efficient design. For some values of $r_{(1)}$, we note $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d_{2}}$ is quite close to $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{*}}$. The value of $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{*}}$ is found to be at most $11\%$, while most of the values of $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d_{2}}$ are much larger. Thus we can conclude that here $d^{*}$ is the best design among the three designs and also is a highly efficient design.
(a) $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{1},d^{*}\}$ (Case $1$)
(b) $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{1},d^{*}\}$ (Case $1$)
(c) $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{1},d^{*}\}$ (Case $2$)
(d) $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{1},d^{*}\}$ (Case $2$)
(e) $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{1},d^{*}\}$ (Case $3$)
(f) $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{1},d^{*}\}$ (Case $3$)
(g) $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{1},d^{*}\}$ (Case $4$)
(h) $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{1},d^{*}\}$ (Case $4$)
Figure 2: Plotting $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ and $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{1},d^{*}\}$, where $p=t=3$, $n=6$ and $\sigma_{11}=\sigma_{22}$.
@@
(Case $4$)
Figure 3: Plotting $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ and $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{1},d_{2},d^{*}\}$, where $p=t=4$, $n=12$ and $\sigma_{11}=\sigma_{22}$.

### v2

@@ 3.2.1 Study of design efficiency for the Markovian covariance structure
$\bm{V}_{R}^{*}$ $=\bm{V}_{R}^{-1}-\left(\bm{{1}}_{p}^{{}^{\prime}}\bm{V}_{R}^{-1}\bm{{1}}_{p}\right)^{-1}\bm{V}_{R}^{-1}\bm{J}_{p\times p}\bm{V}_{R}^{-1}$
$=\frac{1}{r_{(1)}^{4}-4r_{(1)}^{2}+3}\begin{bmatrix}\frac{2\left(r_{(1)}^{4}-4r_{(1)}^{2}+3\right)}{r_{(1)}^{6}-3r_{(1)}^{4}-r_{(1)}^{2}+3}&-1&\frac{r_{(1)}^{4}-4r_{(1)}^{2}+3}{r_{(1)}^{4}-2r_{(1)}^{2}-3}\\ -1&2&-1\\ \frac{r_{(1)}^{4}-4r_{(1)}^{2}+3}{r_{(1)}^{4}-2r_{(1)}^{2}-3}&-1&\frac{2\left(r_{(1)}^{4}-4r_{(1)}^{2}+3\right)}{r_{(1)}^{6}-3r_{(1)}^{4}-r_{(1)}^{2}+3}\end{bmatrix},$
where $\bm{V}_{1}=\sigma_{11}\begin{bmatrix}1&r_{(1)}&r_{(1)}^{2}\\ r_{(1)}&1&r_{(1)}\\ r_{(1)}^{2}&r_{(1)}&1\end{bmatrix}$ and $\bm{V}_{R}=\begin{bmatrix}1&r_{(1)}^{2}&r_{(1)}^{4}\\ r_{(1)}^{2}&1&r_{(1)}^{2}\\ r_{(1)}^{4}&r_{(1)}^{2}&1\end{bmatrix}$. Here, $\sigma_{11}>0$ and $10<r_{(1)}<$ 1. The constraint, $r_{(1)}<1$, is required for the matrix $\bm{V}_{1}$ to be positive definite. To check the necessary and sufficient condition from Remark 3.3, we evaluate $tr\left(\bm{V}_{1}^{*}\bm{\psi}\right)=-\frac{2\sigma_{11}^{-1}}{r_{(1)}^{2}-4r_{(1)}+3}$, $tr\left(\bm{V}_{R}^{*}\bm{\psi}\right)=-\frac{2}{r_{(1)}^{4}-4r_{(1)}^{2}+3}$, $tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{1}^{*}\bm{\psi}\right)=\frac{2\sigma_{11}^{-1}\left(3r_{(1)}+5\right)}{3\left(r_{(1)}^{3}-3r_{(1)}^{2}-r_{(1)}+3\right)}$ and $tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)=$
$\frac{2\left(3r_{(1)}^{2}+5\right)}{3\left(r_{(1)}^{6}-3r_{(1)}^{4}-r_{(1)}^{2}+3\right)}$. Note, these expressions of $\bm{V}_{1}^{*}$, $\bm{V}_{R}^{*}$ and the traces have been cross checked by using MATLAB (The MathWorks Inc., 2024). Since $tr\left(\bm{V}_{1}^{*}\bm{\psi}\right)tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)$ and $tr\left(\bm{V}_{R}^{*}\bm{\psi}\right)\times$
$tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{1}^{*}\bm{\psi}\right)$ are unequal, we may conclude that for $10<r_{(1)}<$ 1, $d^{*}$ does not attain the upper bound of $tr\left(\bm{{C}}_{d(s2)}\right)$ over $d\in\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$. We find that similar results hold for the other cases when $p=t=3$, and for all the cases when $p=t=4$. Thus we may state that for $p=t\in\{3,4\}$, and all cases of $\bm{V}_{1}$ and $\bm{V}_{R}$ considered, $d^{*}$ fails to attain the upper bound $u$.
@@
Here, $RD_{d^{(0)}}$ is a function of $t$, $n$, $p$, $\sigma_{11}$, $\sigma_{22}$, $\rho$, $\bm{V}_{1}$ and $\bm{V}_{R}$, and its values lie between 00 and $1$. For $RD_{d^{(0)}}=0$, the upper bound is attained by the design $d^{(0)}$, thus proving $d^{(0)}$ is an efficient design. Hence, lower values of $RD_{d^{(0)}}$ are preferred as they indicate that $tr\left(\bm{{C}}_{d^{(0)}(s2)}\right)$ is closer to the upper bound $u$, and $d^{(0)}$ is a highly efficient design.
@@
The balanced uniform design $d_{2}$ is constructed using the method described by Bose and Dey, (2009). For $p=t=3$, a balanced uniform design with the same number of subjects as $d^{*}$ is also an orthogonal array, so it is not considered as a candidate for comparison with $d^{*}$. A graphical procedure is used to evaluate the designs $d^{(0)}\in\{d_{1},d_{2},d^{*}\}$. In our illustrations we assume $\sigma_{11}=\sigma_{22}$. The next remark shows that if $\sigma_{11}=\sigma_{22}$, then for $p=t\geq 3$, the relative difference does not depend on the value of $\sigma_{11}$.
@@
From our choices of $\bm{V}_{1}$ and $\bm{V}_{R}$, it is clear that a $p\times p$ matrix $\bm{V}_{1}$ is a function of $\sigma_{11}$ and $r_{(1)}$, while a $p\times p$ matrix $\bm{V}_{R}$ is only a function of $r_{(1)}$. Thus from Remark 3.4, we can see that for any $d^{(0)}\in\mathcal{D}^{(1)}_{t,n=\lambda t\left(t-1\right),p=t}$, $p=t\geq 3$ and $\sigma_{11}=\sigma_{22}$, the relative difference given as $RD_{d^{(0)}}$ is a function of $d^{(0)}$, $p$, $n$, $\rho$ and $r_{(1)}$.
@@
For $p=t\in\{3,4\}$, $n=t\left(t-1\right)$, $\sigma_{11}=\sigma_{22}$, and $d^{(0)}\in\{d_{1},d_{2},d^{*}\}$, we plot $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ and $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$, corresponding to various positive permissible values of $r_{(1)}$. Here by permissible values, we mean those values for which the matrices $\bm{V}_{1}$ and $\bm{V}_{R}$ are positive definite. If in any case $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ attains 00, then this implies that design $d^{(0)}$ maximizes $tr\left(\bm{{C}}_{d^{(0)}(s2)}\right)$ and $d^{(0)}$ is a efficient design. If $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ does not attain 00, we plot $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ to get an understanding of the worst scenario and quantify the departure of the design under consideration from an efficient design. Lower values of $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ implies better performance of the design in terms of efficiency. For $d^{(0)}\in\{d_{1},d^{*}\}$, where $p=t=3$ and $n=6$, Fig. 4 and Fig. 5 display the plots of $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ and $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$, for cases $1$-$7$. While, for $p=t=4$, $n=12$, Fig. 6 and Fig. 7 show the design performances.
@@
From Fig. 4 and Fig. 5, for $p=t=3$ and $n=6$, we see that $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d_{1}}$ never attains 00 and hence $d_{1}$, like $d^{*}$, fails to attain the upper bound $u$. Also, we note that for cases $1$-$7$, the values of $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d_{1}}>>max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{*}}$, implying that $d^{*}$ performs much better when compared to $d_{1}$. Considering all seven cases, the maximum value of $RD_{d^{*}}$ over $\rho\neq 0$ is at most around $0.39\%$ indicating that $d^{*}$ is a highly efficient design. In the $p=t=4$ and $n=12$ case, we observe from Fig. 6 and Fig. 7 that for none of the designs $d^{(0)}\in\{d_{1},d_{2},d^{*}\}$, $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ attain 00 and $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{*}}<max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d_{2}}<max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d_{1}}$. So $d^{*}$ performs the best among the three designs being compared, followed by $d_{2}$. Since $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d_{1}}$ is far away from 00, we reject the candidature of $d_{1}$ as a highly efficient design. For some values of $r_{(1)}$, we note $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d_{2}}$ is quite close to $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{*}}$. The value of $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{*}}$ is found to be at most $2.02\%$, while most of the values of $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d_{2}}$ are much larger. Thus, we may conclude that $d^{*}$ is the best design among the three designs being compared and also is a highly efficient design.
(a) $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{1},d^{*}\}$
($\bm{V}_{1}$: Mat($0.5$), $\bm{V}_{R}$: Mat($1.5$))
(b) $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{1},d^{*}\}$
($\bm{V}_{1}$: Mat($0.5$), $\bm{V}_{R}$: Mat($1.5$))
(c) $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{1},d^{*}\}$
($\bm{V}_{1}$: Mat($0.5$), $\bm{V}_{R}$: Mat($\infty$))
(d) $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{1},d^{*}\}$
($\bm{V}_{1}$: Mat($0.5$), $\bm{V}_{R}$: Mat($\infty$))
(e) $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{1},d^{*}\}$
($\bm{V}_{1}$: Mat($1.5$), $\bm{V}_{R}$: Mat($0.5$))
(f) $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{1},d^{*}\}$
($\bm{V}_{1}$: Mat($1.5$), $\bm{V}_{R}$: Mat($0.5$))
(g) $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{1},d^{*}\}$
($\bm{V}_{1}$: Mat($1.5$), $\bm{V}_{R}$: Mat($\infty$))
(h) $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{1},d^{*}\}$
($\bm{V}_{1}$: Mat($1.5$), $\bm{V}_{R}$: Mat($\infty$))
Figure 4: Plots of $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ and $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{1},d^{*}\}$, $p=t=3$, $n=6$ and $\sigma_{11}=\sigma_{22}$.
(a) $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{1},d^{*}\}$
($\bm{V}_{1}$: Mat($\infty$), $\bm{V}_{R}$: Mat($0.5$))
(b) $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{1},d^{*}\}$
($\bm{V}_{1}$: Mat($\infty$), $\bm{V}_{R}$: Mat($0.5$))
(c) $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{1},d^{*}\}$
($\bm{V}_{1}$: Mat($\infty$), $\bm{V}_{R}$: Mat($1.5$))
(d) $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{1},d^{*}\}$
($\bm{V}_{1}$: Mat($\infty$), $\bm{V}_{R}$: Mat($1.5$))
(e) $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{1},d^{*}\}$
($\bm{V}_{1}$: Mat($0.5$), $\bm{V}_{R}$: Mat($0.5$))
(f) $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{1},d^{*}\}$
($\bm{V}_{1}$: Mat($0.5$), $\bm{V}_{R}$: Mat($0.5$))
Figure 5: Plots of $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ and $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{1},d^{*}\}$, $p=t=3$, $n=6$ and $\sigma_{11}=\sigma_{22}$.
@@
($\bm{V}_{1}$: Mat($1.5$), $\bm{V}_{R}$: Mat($\infty$))
Figure 6: Plots of $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ and $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{1},d_{2},d^{*}\}$, $p=t=4$, $n=12$ and $\sigma_{11}=\sigma_{22}$.
($\bm{V}_{1}$: Mat($\infty$), $\bm{V}_{R}$: Mat($0.5$))
($\bm{V}_{1}$: Mat($\infty$), $\bm{V}_{R}$: Mat($0.5$))
($\bm{V}_{1}$: Mat($\infty$), $\bm{V}_{R}$: Mat($1.5$))
($\bm{V}_{1}$: Mat($\infty$), $\bm{V}_{R}$: Mat($1.5$))
($\bm{V}_{1}$: Mat($0.5$), $\bm{V}_{R}$: Mat($0.5$))
($\bm{V}_{1}$: Mat($0.5$), $\bm{V}_{R}$: Mat($0.5$))
Figure 7: Plots of $min_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ and $max_{0<|\rho|<1}\leavevmode\nobreak\ RD_{d^{(0)}}$ for $d^{(0)}\in\{d_{1},d_{2},d^{*}\}$, $p=t=4$, $n=12$ and $\sigma_{11}=\sigma_{22}$.